OpenStudy (anonymous):

@Mayaal What is the Euler characteristic of a polyhedron that has 6 vertices, 8 faces, and 12 edges?

OpenStudy (anonymous):

I used F+V-E=2

OpenStudy (anonymous):

Hello again :D

OpenStudy (mayaal):

hi

OpenStudy (anonymous):

Ok i got 2 atleast i think O:

OpenStudy (anonymous):

The option for the question are -2 2 14 10

OpenStudy (mayaal):

Let me think...

OpenStudy (anonymous):

I did this V+F-E=2 6+8-12=2 14-12=2 2=2

OpenStudy (mayaal):

do you something about Eulers characteristics?

OpenStudy (mayaal):

know*

OpenStudy (anonymous):

Vaguely

OpenStudy (mayaal):

let me know...

OpenStudy (anonymous):

I know we are going to go into his theorems sooon

OpenStudy (anonymous):

Will do!

OpenStudy (mayaal):

I think the work u did is absolutely right...!

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

I went to wikipedia to refresh

OpenStudy (anonymous):

His charestics are the formulas you showed me!

OpenStudy (mayaal):

:) CoNgRaTs you did it correctly

OpenStudy (anonymous):

Since i used V+F-E=2

OpenStudy (mayaal):

Yup so its correct

OpenStudy (mayaal):

No need for a medal you did it all the work by yourself